3 consecutive odd intergers whose sum is -57.
Answer by fractalier(2101)

Let the numbers be x, x+2, and x+4.
Thus x + x+2 + x+4 = -57 and
3x + 6 = -57
3x = -63
x = -21
x+2 = -19
x+4 = -17